OpenStudy (anonymous):

If $600 is invested at an interest rate of 2.5% per year, find the amount of the investment at the end of 10 years for the following compounding methods: a) annually b) semiannually c) quarterly d) continuously

OpenStudy (anonymous):

\[600\times (1.025)^{10}\] \[600\times (1+\frac{025}{2})^{20}\] \[600\times (1+\frac{025}{4})^{40}\] \[600\times e^{.25}\]

OpenStudy (anonymous):

typo there middle two are \[600\times (1+\frac{.025}{2})^{20}\] \[600\times (1+\frac{.025}{4})^{40}\]

OpenStudy (anonymous):

calculator for the result

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

here is the third one http://www.wolframalpha.com/input/?i=600%281%2B.025%2F4%29^40

OpenStudy (anonymous):

second one http://www.wolframalpha.com/input/?i=600%281%2B.025%2F2%29^20

OpenStudy (anonymous):

fourth one http://www.wolframalpha.com/input/?i=600e^.25 you can see that they are close